John invested $11000, a portion earning a simple interest rate of 4.5% per year and the rest earning a simple intrest rate of 4% per year. After one year the total value of John's investment is $11445.50.
How much did John invest at 4.5% interest rate? And how much did John invest at 4% interest rate?
HINT: the total value of John's investment includes both the amount he invested and the interest he earned.

Answer by stanbon(48558)

John invested $11000, a portion earning a simple interest rate of 4.5% per year and the rest earning a simple intrest rate of 4% per year. After one year the total value of John's investment is $11445.50.
How much did John invest at 4.5% interest rate? And how much did John invest at 4% interest rate?
--------------------
Equation:
interest + interest = interest
0.045x + 0.04(11000-x) = 445.50
Multiply thru by 1000 to get:
45x + 40*11000 - 40x = 445500
5x = 5500
x = $1100 (amt invested at 4.5%)
11000 - 1100 = $9900 (amt. invested at 4%
============================================
Cheers,
Stan H.